# Gröbner Basis Cryptanalysis of Ciminion and Hydra

### v1

@@ 7.1 Ciminion
$r_{C}$ $r_{E}$ Fast Polynomial Multiplication Eigenvalue Method DRL Computation Iterated Model DRL Computation Fully Substituted Model [DGGK21, §4.4]
$30$ $6$ $49$ $68$ $134$ $72$
$60$ $6$ $80$ $128$ $253$ $132$
$90$ $6$ $111$ $188$ $372$ $192$
$90$ $14$ $119$ $204$ $404$ $208$
$90$ $23$ $128$ $222$ $440$ $226$
Also, recall that $r_{C}\in\left\{s+6,\left\lceil\frac{2\cdot(s+6)}{3}\right\rceil\right\}$, where $s$ is the security level, for two Ciminion instantiations, see Table 2. In particular, for Ciminion2 this already implies that for the quotient space dimension $\dim_{\mathbb{F}_{q}}\left(\mathcal{F}_{\texttt{Ciminion}}\right)\geq 2^{\frac{s}{2}}$. Therefore, round numbers based on Table 2 should be sufficient to achieve the security level in Gröbner basis attacks via the iterated polynomial system, provided of course a fix against Bariant’s attack [Bar23] is implemented.
7.2 Hydra
For Gröbner basis cryptanalysis of Hydra the designers considered the downsized iterated Hydra polynomial system, see [GØSW23, §7.4] and [GØSW22, Appendix I.2]. The designers assumed that the downsized system of $22\cdot r_{\mathcal{H}}+$ 2 equations in $22\cdot r_{\mathcal{H}}-$ 2 variables is semi-regular [Frö85, Par10]. Under the semi-regularity assumption one can extract the degree of regularity ([BFS04, Definition 4]) $d_{\operatorname{reg}}$ of the Hydra polynomial system from the Hilbert series [KR05, Chapter 5] which then yields the complexity estimate ([BFS04, Theorem 7])
$\mathcal{O}\left((2\cdot r_{\mathcal{H}}+2)\cdot d_{\operatorname{reg}}\cdot\binom{2\cdot r_{\mathcal{H}}-2+d_{\operatorname{reg}}}{d_{\operatorname{reg}}}^{\omega}\right)$ (76)
for DRL Gröbner basis computations. For $r_{\mathcal{H}}^{\ast}=29$ the designers estimate $D_{\operatorname{reg}}=23$, with $\omega=2$ this yields a complexity estimate of $\approx 2^{130.8}$ bits. Moreover, it is claimed that $r_{\mathcal{H}}=r_{\mathcal{H}}^{\ast}+2=31$ rounds are sufficient to provide $128$ bits of security against Gröbner basis attacks.
7.2.1 Term Order Conversion & Eigenvalue Method Complexity Estimate
In Section 4.2 we have outlined how one can extract a zero-dimensional DRL Gröbner basis from the Hydra polynomial system via a linear change of coordinates. We can directly use this Gröbner basis to perform term order conversion to LEX, or apply the dedicated Eigenvalue Method from Section 5 combined with $\mathbb{F}_{q}$-valued root extraction via the GCD. Since we could not prove that a LEX Hydra Gröbner basis is in shape position, we have to use the complexity for the probabilistic FGLM [FGHR14] of radical ideals which becomes
$\tilde{\mathcal{O}}\left((r_{\mathcal{H}}-1)\cdot 2^{\omega\cdot 2\cdot(r_{\mathcal{H}}-1)+1}\right).$ (77)
For the Eigenvalue Method, the Hydra DRL Gröbner basis satisfies the assumptions of Section 5, and in addition it satisfies the shape condition of Equation 58. Therefore, we can use the complexity estimate from Equation 61. Since we do not know the number of $\mathbb{F}_{q}$-valued eigenvalues for the Hydra multiplication matrices we opt for a better safe than sorry approach. We assume that the first eigenvalue that an adversary recovers for each variable leads to the true key, i.e. we assume $N=1$. Then, the Eigenvalue Method complexity becomes
$\mathcal{O}\left(2^{\omega}\cdot\frac{2^{\omega\cdot(2\cdot r_{\mathcal{H}}-2)}-1}{2^{2\cdot\omega}-1}+\sum_{i=1}^{r_{\mathcal{H}}-1}D\big{(}2^{2\cdot i},q\big{)}\right),$ (78)
where $D\big{(}2^{i},q\big{)}$ again denotes the complexity of the GCD with the field equation (Equation 1).
In Table 4 we have computed the complexities for term order conversion and efficient $\mathbb{F}_{q}$-valued root extraction via the GCD. From the formulae, we can directly see that for $\omega=2$ and $r_{\mathcal{H}}\geq 34$ the complexities will always exceed $128$ bits. On the other, hand for $r_{\mathcal{H}}^{\ast}=29$ term order conversion requires only $118$ bits, and for $r_{\mathcal{H}}=r_{\mathcal{H}}^{\ast}+2=31$ it requires $126$ bits. This invalidates the claim by the Hydra designers that $31$ rounds are sufficient to protect against Gröbner basis attacks. Moreover, term order conversion always outperforms DRL Gröbner basis computation under the semi-regularity assumption. Lastly, for the round numbers in Table 4 the Eigenvalue Method shaves-off at least $6$ additional bits compared to term order conversion.
Table 4: Complexity estimate of the Eigenvalue Method and term order conversion to LEX with efficient $\mathbb{F}_{q}$-valued root extraction for Hydra with $q=2^{127}+45$ and $\omega=2$.
@@ 7.2.1 Term Order Conversion & Eigenvalue Method Complexity Estimate
$r_{\mathcal{H}}$ Eigenvalue Method Term Order Conversion Semi-Regular Estimate [GØSW23]
$20$ $75$ $82$ $89$
$29$ $111$ $118$ $131$
$31$ $119$ $126$ $142$
$33$ $127$ $135$ $153$
$34$ $131$ $139$ $159$
$39$ $151$ $159$ $183$
7.2.2 Proven DRL Complexity Estimate
Via Theorem 4.1 and our SageMath implementation (Section 4.3) we obtain the proven DRL solving degree estimate for the Hydra polynomial instance from Section 4.3.1
$\operatorname{sd}_{DRL}\left(\mathcal{F}_{\textsf{Hydra}}\right)\leq 2\cdot r_{\mathcal{H}}.$ (79)
Since the Hydra polynomial system contains a zero-dimensional DRL Gröbner basis after a linear change of coordinates (Section 4.2), we can reduce Gaussian elimination on the full Macaulay matrix to Gaussian elimination on the $\mathbb{F}_{q}$-Boolean Macaulay matrix. With Theorem 6.4 and Equation 65 the complexity of constructing the $\mathbb{F}_{q}$-Boolean Macaulay matrix and Gaussian elimination for Hydra becomes
$\mathcal{O}\left(\left(8\cdot r_{\mathcal{H}}^{2}-4\cdot r_{\mathcal{H}}\right)\cdot\left(\sum_{i=0}^{2\cdot r_{\mathcal{H}}-2}\binom{2\cdot r_{\mathcal{H}}+i-3}{i}\cdot\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)+2^{\omega\cdot 2\cdot(r_{\mathcal{H}}-1)+2}\right).$ (80)
In Table 5 we evaluate our complexity bound.
Table 5: DRL Gröbner basis complexity estimate for Hydra via Equation 80 with $\omega=2$.
Complexity (Bits)
$r_{\mathcal{H}}$ Proven Estimate Semi-Regular Estimate [GØSW23]
$20$ $105$ $89$
$29$ $151$ $131$
$31$ $161$ $142$
$33$ $172$ $153$
$34$ $177$ $159$
$39$ $203$ $183$
Therefore, we have derived a computer aided proof for the complexity of Gröbner basis attacks on Hydra without resorting to unproven assumptions like semi-regularity. Our proven complexity estimate is significantly larger than the hypothesized one for Hydra, since the degree of regularity estimate is always less than the Macaulay bound.
7.2.3 Semi-Regular Complexity Estimate for the Boolean Macaulay Matrices
As already mentioned, in the Hydra cryptanalysis it is assumed that the polynomial system of $22\cdot r_{\mathcal{H}}+$ 2 equations in $22\cdot r_{\mathcal{H}}-$ 2 variables is semi-regular. If we assume that the substituted Hydra polynomial system that contains the DRL Gröbner basis (Section 4.2) is also semi-regular, then we can again improve from the generic semi-regular complexity estimate (Equation 76) by passing to the $\mathbb{F}_{q}$-Boolean Macaulay matrix. Again, we can then use Equation 65 to obtain the complexity of construction and Gaussian elimination on the $\mathbb{F}_{q}$-Boolean Macaulay matrices
$\mathcal{O}\left(\begin{aligned} \left(8\cdot r_{\mathcal{H}}^{2}-4\cdot r_{\mathcal{H}}\right)&\cdot\left(\sum_{i=0}^{d_{\operatorname{reg}}}\binom{2\cdot r_{\mathcal{H}}+i-3}{i}\cdot\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)\\ &+4\cdot\left(\sum_{i=0}^{d_{\operatorname{reg}}}\binom{2\cdot r_{\mathcal{H}}-2}{i}\right)^{\omega}\end{aligned}\right).$ (81)
In Table 6 we evaluate this complexity estimate. In addition, we show the complexity just for Gaussian elimination on the $\mathbb{F}_{q}$-Boolean Macaulay matrices, i.e. the complexity under the assumption that the Boolean Macaulay matrices are constructed “for free”. If we account for the cost of constructing the Boolean Macaulay matrices, then this method performs worse than the optimized Eigenvalue Method. However, it still beats the cryptanalysis of the Hydra designers for larger round numbers. Though, if we just consider the Gaussian elimination step, then this method slightly improves upon the Eigenvalue method.
Table 6: DRL Gröbner basis complexity estimate for Hydra via Equation 80 with $\omega=2$.
Complexity (Bits)
$r_{\mathcal{H}}$ $d_{\operatorname{reg}}$ Boolean Gaussian Elimination Boolean Semi-Regular Estimate Semi-Regular Estimate [GØSW23]
$20$ $16$ $74$ $91$ $89$
$29$ $23$ $108$ $130$ $131$
$31$ $25$ $116$ $140$ $142$
$33$ $27$ $125$ $149$ $153$
$34$ $28$ $129$ $154$ $159$
$39$ $32$ $148$ $176$ $183$
Empirical F4 Working Degree.
The Hydra designers performed small scale Gröbner basis experiments for Hydra with the F4 [Fau99] algorithm, see [GØSW22, Table 4]. We replicated their experiment with our OSCAR implementation over the prime field $\mathbb{F}_{7741}$. We report our results in Table 7, but as input system we use the substituted system from Section 4.2 which contains a zero-dimensional DRL Gröbner basis. Our experiments have been performed on an AMD EPYC-Rome (48) CPU with 94 GB RAM. Like the Hydra designers we observed that the empirical F4 working degree meets the semi-regular degree of regularity estimate.
Table 7: Highest working degree and running times for F4 in small scale Hydra polynomial systems over the prime field $\mathbb{F}_{7741}$.
Our data Grassi et al.’s data [GØSW22, Table 4]
$r_{\mathcal{H}}$ $3$ $4$ $5$ $6$ $7$ $8$ $9$ $10$ $6$ $7$ $8$ $9$ $10$ $11$
Highest F4 working degree $3$ $3$ $4$ $5$ $5$ $6$ $7$ $8$ $5$ $5$ $6$ $7$ $8$ $8$
Total time (s) $0.01$ $0.01$ $0.01$ $0.03$ $0.21$ $2.65$ $66.85$ $2295.7$ $0.1$ $0.9$ $13.4$ $293.8$ $8799.5$ $229606.8$
Overall, we conclude that currently the most competitive algebraic attacks against Hydra directly apply polynomial system solving techniques to the zero-dimensional DRL Gröbner basis, see Table 4.
@@ Acknowledgments
Matthias Steiner has been supported in part by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 725042) and the enCRYPTON project (grant agreement No. 101079319).

### v4

@@ 4.4.2 Recomputing the DRL Gröbner Basis
Proven Solving Degree Semi-Regular
$r_{\mathcal{H}}$ $d_{\operatorname{reg}}$ Boolean Matrix Construction Boolean Gaussian Elimination Boolean Matrix Construction Boolean Gaussian Elimination Semi-Regular Estimate [GØSW23]
$28$ $22$ $148.72$ $110$ $118.28$ $102.10$ $125.31$
$29$ $23$ $153.89$ $114$ $123.29$ $106.25$ $130.80$
$30$ $24$ $159.05$ $118$ $128.27$ $110.40$ $136.29$
$31$ $25$ $164.21$ $122$ $133.25$ $114.54$ $141.77$
$32$ $26$ $169.37$ $126$ $138.21$ $118.66$ $147.26$
$33$ $27$ $174.52$ $130$ $143.16$ $122.79$ $152.75$
$34$ $28$ $179.67$ $134$ $148.10$ $126.90$ $158.23$
$35$ $28$ $184.82$ $138$ $150.37$ $129.66$ $160.24$
$39$ $32$ $205.41$ $154$ $170.12$ $146.16$ $182.22$
$45$ $37$ $236.24$ $178$ $197.03$ $169.55$ $211.72$
Empirical F4 Working Degree.
The Hydra designers performed small scale Gröbner basis experiments for Hydra with the F4 [Fau99] algorithm, see [GØSW22, Table 4]. With our OSCAR implementation we replicated these experiments to verify that F4’s highest working degree matches the semi-regular degree of regularity before and after the change of coordinates. Our results are recorded in Table 5. Our experiments have been performed on an AMD EPYC-Rome (48) CPU with 94 GB RAM.
In all our experiments F4’s highest working degree indeed matches the semi-regular degree of regularity. As expected, after the change of coordinates the computation is faster, because F4 will implicitly eliminate all monomials which are divisible by a square.
Table 5: Hydra highest working degree and total running time of F4 over the prime field $\mathbb{F}_{7741}$. The column $d_{\operatorname{reg}}$ shows the degree of regularity under the semi-regularity assumption.
Before Change of Coordinates After Change of Coordinates Grassi et al.’s data [GØSW22, Table 4]
$r_{\mathcal{H}}$ $d_{\operatorname{reg}}$ Highest F4 working degree Total time (s) Highest F4 working degree Total time (s) Highest F4 working degree Total time (s)
$3$ $3$ $3$ $0.01$ $3$ $0.01$
$4$ $3$ $3$ $0.01$ $3$ $0.01$
$5$ $4$ $4$ $0.02$ $4$ $0.01$
$6$ $5$ $5$ $0.16$ $5$ $0.03$ $5$ $0.1$
$7$ $5$ $5$ $1.35$ $5$ $0.25$ $5$ $0.9$
$8$ $6$ $6$ $14.88$ $6$ $2.68$ $6$ $13.4$
$9$ $7$ $7$ $384.13$ $7$ $67.19$ $7$ $293.8$
$10$ $8$ $8$ $8676.30$ $8$ $3194.95$ $8$ $8799.5$
$11$ $8$ $8$ $229606.8$
4.4.3 Term Order Conversion & Eigenvalue Method
With a Hydra DRL Gröbner basis at hand, see Section 4.2, we can proceed to direct system solving techniques. For example, we can use a generic algorithm like the state-of-the-art probabilistic FGLM algorithm [FGHR14] to convert to a LEX Gröbner basis, and then factor the univariate polynomial. Assuming that the Hydra DRL Gröbner basis is in shape position the complexity of term order conversion is bounded by [FGHR14, Proposition 3]
$\mathcal{O}\Big{(}(2\cdot r_{\mathcal{H}}-2)\cdot 2^{2\cdot r_{\mathcal{H}}-2}\cdot\big{(}2^{(\omega-1)\cdot(2\cdot r_{\mathcal{H}}-2)}+(2\cdot r_{\mathcal{H}}-2)^{2}\cdot\log_{2}(2\cdot r_{\mathcal{H}}-2)\big{)}\Big{)}$ (64)
field operations. Next we have to factor the univariate polynomial of degree $2^{2\cdot r_{\mathcal{H}}-2}$ via the GCD with the field equation. However, even for full rounds $r_{\mathcal{H}}=39$ this degree is well below $2^{128}$. Therefore, the GCD complexity from Equation 21 will not exceed $128$ bits, and henceforth can be ignored.
On the other hand, the Hydra DRL Gröbner basis satisfies the structure assumption from Equation 22, so we can pass to the dedicated Eigenvalue Method to find a $\mathbb{F}_{q}$-valued solution. For Hydra, the complexity from Equation 25 then evaluates to
$\mathcal{O}\left(2^{\omega+1}\cdot\frac{2^{\omega\cdot(2\cdot r_{\mathcal{H}}-2)}-N^{r_{\mathcal{H}}-1}}{2^{2\cdot\omega}-N}\right),$ (65)
where $N$ is a bound on the number of $\mathbb{F}_{q}$-valued solutions in every iteration. Again, the GCD complexity can be considered negligible. Since a bound on $N$ is in general not available, we consider a worst-case scenario approach. We assume that the first root which an adversary recovers in every iteration leads to the true Hydra key, i.e. we use $N=1$ in Equation 65.
In Table 6 we evaluated the complexities for direct system solving techniques and compared them to the other strategies. Our estimations are again from a designer’s point of view which assumes an ideal adversary who can achieve the linear algebra constant $\omega=2$. In this scenario, term order conversion breaks up to $r_{\mathcal{H}}=31$ rounds below $128$ bits. This invalidates the claim of the Hydra designers [GØSW23, §7.4] that $r_{\mathcal{H}}=31$ rounds are sufficient to provide $128$ bits of security against Gröbner basis attacks for an ideal adversary. Moreover, via the dedicated Eigenvalue Method up to $r_{\mathcal{H}}=33$ do not achieve $128$ bits of security. For full rounds $r_{\mathcal{H}}=39$ the security margin is reduced by $\approx 30$ bits, however the security claim is not affected.
Let $r_{\mathcal{H}}^{\ast}$ be the minimum round number which achieves $128$ bits of security in a Gröbner basis attack, i.e. according to our analysis we have $r_{\mathcal{H}}^{\ast}=34$. Recall from Equation 12 that the head round number is set as $r_{\mathcal{H}}=\left\lceil 1.25\cdot\max\left\{24,2+r_{\mathcal{H}}^{\ast}\right\}\right\rceil$. Thus, our analysis suggests that this round number is increased to $r_{\mathcal{H}}=45$ to recover the security margin originally intended by the designers.
In case one wants to be extra conservative, one can consider the construction of the $\mathbb{F}_{q}$-Boolean Macaulay matrix to be for free. Then, we have $r_{\mathcal{H}}^{\ast}=35$, see Table 4, which implies $r_{\mathcal{H}}=47$.
Table 6: Hydra complexity estimations for term order conversion and the Eigenvalue Method. The column $d_{\operatorname{reg}}$ shows the degree of regularity under the semi-regularity assumption. All estimations use $\omega=2$.
@@ 4.4.3 Term Order Conversion & Eigenvalue Method
$r_{\mathcal{H}}$ $d_{\operatorname{reg}}$ Term Order Conversion Eigenvalue Method Boolean Semi-Regular Estimate Semi-Regular Estimate [GØSW23]
$28$ $22$ $113.75$ $107.09$ $118.28$ $125.31$
$29$ $23$ $117.81$ $111.09$ $123.29$ $130.80$
$30$ $24$ $121.86$ $115.09$ $128.27$ $136.29$
$31$ $25$ $125.91$ $119.09$ $133.25$ $141.77$
$32$ $26$ $129.95$ $123.09$ $138.21$ $147.26$
$33$ $27$ $134.00$ $127.09$ $143.16$ $152.75$
$34$ $28$ $138.04$ $131.09$ $148.10$ $158.23$
$35$ $28$ $142.09$ $135.09$ $150.37$ $160.24$
$39$ $32$ $158.25$ $151.09$ $170.12$ $182.22$
$45$ $37$ $182.46$ $175.09$ $197.03$ $211.72$
Small Scale Experiments.
With our OSCAR implementation we computed the characteristic polynomial of the Hydra multiplication matrix $\mathbf{M}_{\hat{x}_{2\cdot r_{\mathcal{H}}-2}}$, see Equation 19, for small $r_{\mathcal{H}}$ and measured the time of the overall computation. In addition, we computed the polynomial matrix determinant from Equation 19 for uniformly random matrices $\mathbf{A}_{0},\mathbf{A}_{1}\in\mathbb{F}_{q}^{(2\cdot r_{\mathcal{H}}-3)\times(2\cdot r_{\mathcal{H}}-3)}$. For construction and determinant of the matrix we used OSCAR’s functions for multivariate polynomial division and determinant respectively. Our results are recorded in Table 7. In addition, we performed term order conversion to LEX for our Hydra DRL Gröbner basis with OSCAR’s fglm function. For comparison, we also generated a random multivariate quadratic polynomial system in the same number of variables as Hydra, computed the DRL Gröbner basis, and finally converted to LEX via FGLM. Our FGLM results are recorded in Table 8. All experiments in this section have been performed on an AMD EPYC-Rome (48) CPU with 94 GB RAM.
For the characteristic polynomial, we observed that the construction time for the multiplication matrix is negligible compared to the one for the characteristic polynomial. Moreover, Hydra characteristic polynomials are computed faster than for random matrices $\mathbf{A}_{0}$ and $\mathbf{A}_{1}$.
For term order conversion, we observed that FGLM on Hydra systems requires less time than for random multivariate quadratic systems. We also note that in all our experiments the Hydra LEX Gröbner basis was in $\hat{x}_{2\cdot r_{\mathcal{H}}-2}$-shape position.
Table 7: Hydra running times for construction of the multiplication matrix and its characteristic polynomial over the prime field $\mathbb{F}_{7741}$. The column “Determinant Random Matrix” shows the running time of the determinant from Equation 19 with uniformly random $\mathbf{A}_{0},\mathbf{A}_{1}\in\mathbb{F}_{7741}^{(2\cdot r_{\mathcal{H}}-3)\times(2\cdot r_{\mathcal{H}}-3)}$.
Time (s)
$r_{\mathcal{H}}$ Construction Hydra Matrix Characteristic Polynomial Hydra Matrix Determinant Random Matrix
$5$ $<0.1$ $0.3$ $0.4$
$6$ $0.7$ $73.9$ $82.8$
$7$ $11.1$ $8336.7$ $11722.9$
Table 8: Hydra FGLM running times over the prime field $\mathbb{F}_{7741}$. The column “Random Quadratic F4” shows the running time for computing a DRL Gröbner basis with F4 for a random multivariate quadratic polynomial system in the same number of variables as the Hydra system. The column “Random Quadratic FGLM” shows the running time of FGLM for the random DRL Gröbner basis.
Time (s)
$r_{\mathcal{H}}$ Hydra FGLM Random Quadratic F4 Random Quadratic FGLM
$5$ $0.3$ $0.1$ $0.4$
$6$ $20.0$ $0.8$ $26.1$
$7$ $1265.7$ $15.6$ $1598.5$
$8$ $78345.1$ $1123.9$ $99492.9$
@@ Acknowledgments
Matthias Steiner has been supported in part by the enCRYPTON project (grant agreement No. 101079319).